OpenStudy (anonymous):

show work 6(x-5)-(x+6)

pooja195 (pooja195):

Distribute the Negative Sign: =6(x−5)+−1(x+6) =6(x−5)+−1x+(−1)(6) =6(x−5)+−x+-6

pooja195 (pooja195):

Now distribute the 6 (6)(x)+(6)(−5)+−x+−6 what are you left with?

OpenStudy (anonymous):

6x-6(5)-(x+6) 6x-30-(x+6) 6x-x-30+6 5x-30+6 5x-24

OpenStudy (anonymous):

5x-36?

OpenStudy (anonymous):

sorry its 5x-36 i forgot to change 6 to negative

OpenStudy (anonymous):

thank you